# Rayleigh-Bénard convective motion of stratified fluids in the Earth’s troposphere

### v1

@@ II.1 Convective flow in an unbounded domain
For visualization of the characteristics of the growth rate, we have plotted $\gamma$ against the normalized wave number $k_{\perp}H$ as shown in Fig. 1. It is evident that, in contrast to Cases I and II, where the buoyancy force is the only dominating force and the perturbations grow with increasing values of $k_{\perp}H$, the growth rate has cut-offs at smaller and larger wavelengths of perturbations. First of all, from the dependency of $\gamma$ [See Eq. (16)] on the parameters, we note that the qualitative behaviors of $\gamma$ by the influences of $\beta$ and $\chi$ are the same as of $\Gamma$ and $\nu$ respectively. Thus, it is sufficient to investigate the qualitative features of $\gamma$ by the effects of the parameters $\beta$, $\chi$, $N^{2}$ (for the present case it is assumed to be zero), and the vertical wave number $k_{z}H$. Initially, for fixed values of the parameters $\beta$, $\chi$, and $k_{z}H$, and a small value of $k_{\perp}H~{}(<1)$, the contributions from both the buoyancy and dissipative forces are small and so is the growth rate (close to zero). As the value of $k_{\perp}H$ starts increasing, the buoyancy force tends to dominate over the dissipation for which the growth rate gradually increases and reaches a maximum value until the domination is strong. As $k_{\perp}H$ further increases, the contribution from the dissipation starts increasing and the growth rate tends to reduce and eventually vanishes at a finite value of $k_{\perp}H$ (typically $>1$). Thus, it may be concluded that in a dissipative medium of incompressible fluids under gravity in an unbounded domain of the troposphere, the perturbations can grow (leading to the convective instability) only within a finite domain of the wave number $k_{\perp}H$ due to an interplay between the buoyancy force and the dissipation. On small increasing the value of the thermal expansion coefficient $\beta$, and hence the contribution from the buoyancy force, a significant enhancement of the growth rate having a cut-off at higher $k_{\perp}H$ is seen (See the dashed line). This implies that as the wave number or the vertical length scale of perturbations deepen, the buoyancy force driven wave motion tends become more unstable with higher growth rates. However, such growth rates can be reduced by increasing either $\chi$ (i.e., enhanced contribution from the dissipation) or the vertical wave number $k_{z}H$ (Since the contribution of the buoyancy force inversely varies with $k_{z}H$, it decreases with increasing $k_{z}H$; See the dotted and dash-dotted lines).
@@
Figure 1: The instability growth rate, given by Eq. (16) ($\gamma\equiv\gamma_{+}$), is shown (Case III, Sec. II.1: $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}=0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, and $k_{z}H$ as in the legend. Here, $H$ is the vertical length scale of perturbation. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$.
@@
We note that the lower sign of $\pm$ in $\gamma$ gives a purely damped mode due to the effects of the kinematic viscosity and the thermal diffusivity, which is not of interest to the present study. Thus, for the upper sign, the instability condition $(\gamma>0)$ gives
@@
From Eq. (25), we find that for a given vertical scale height $H$, the term proportional to $\beta$, associated with the buoyancy force due to thermal expansion and background density inhomogeneity, reduces the critical (or minimum) value of the Rayleigh number [cf. Eq. (20)] above which the convective instability can occur. Typically, for $H\sim 800$ m, $N^{2}\sim-0.003/\rm{s}^{2}$, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=260$ K, $\Gamma=0.0065$ K/m, $\beta=0.001$/K, $\chi=3000~{}\rm{m^{2}/s}$, and $\nu=2000~{}\rm{m^{2}/s}$, relevant for tropospheric fluids [6], the variation of the critical values of the Rayleigh number against the normalized wave number $k_{\perp}H$ are shown in Fig. 2. It is seen that after the sharp peak at a point close to the singular point $k_{\perp}H=0$, the critical value of the Rayleigh number decreases achieving minimum at $k_{\perp}H\approx 0.35$, and then tends to increase slowly with increasing values of $k_{\perp}H$. Thus, given a small but finite value of the wave number $k_{\perp}$, a larger vertical length scale of perturbations $H$ results in higher values of $Ra_{\rm{min}}$ and the gravitational force to become more dominant over the pressure gradient and viscous forces, which can lead to stronger convective instability. Mathematically, the minimum value of the Rayleigh number can be exceedingly high at $k_{\perp}H\lesssim 0.1$, i.e., when the wavelength of horizontal perturbations greatly exceeds the vertical scale size $H$. However, this limiting case may not be relevant to an unbounded vertical domain (with infinite extent) of the troposphere. We also note that the critical Rayleigh number assumes values less than unity in the interval $0.2\lesssim k_{\perp}H\lesssim 0.7$, which corresponds to a regime where the dissipative force is stronger than the buoyancy force. In this case, for the instability to occur, the Rayleigh number can be smaller than unity even when it exceeds the critical value. Such a small value of $Ra<1$ and the corresponding regime of $k_{\perp}H$ may not be admissible as we require the buoyancy force to be dominant over the dissipative force for stronger instability. Thus, convective wave motion of tropospheric viscous fluids under gravity becomes more unstable when the wavelengths of horizontal perturbations remain smaller compared to the vertical scale size (i.e., $k_{\perp}^{-1}<H$). We note that the critical value of the Rayleigh number does not alter significantly with a small change of the values of the parameters $\beta$ (or $\Gamma$), $\chi$ (or $\nu$), and $N^{2}$. Also, at $Ra=Ra_{\rm{min}}$, the growth rate vanishes and one can have the convective motion of neutral waves for $N^{2}<0$ with the wave frequency, given by,
@@
Figure 2: The critical value (or minimum value) of the Rayleigh number, given by Eq. (25), is plotted against the perpendicular component of the normalized wave number $k_{\perp}H$. Apart from the sharp peak at a point close to the singular point $k_{\perp}H=0$, the critical value tends to increase slowly with increasing values of $k_{\perp}H$. The fixed parameter values are $\beta=0.001$/K, $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}=0.04$, $k_{z}H=0.5$, and $N^{2}=-0.003/\rm{s}^{2}$.
@@
In particular, for $N^{2}=0$, i.e., when the density inhomogeneity is ignored, the condition (24) reduces to the same condition as Eq. (17).
@@
Figure 3 shows the profiles of the real wave frequency $\omega_{r}$ [Subplot (a)] and the growth rate of instability $\gamma$ [Subplot (b)], given by Eq. (23), for different values of the parameters $\beta$, $\chi$, $N^{2}$, and $k_{z}$. The influences of the other parameters on the growth rate are not shown for the reason mentioned before in Case III. The authors of Ref. [7] discussed the characteristics of $\omega_{r}$ and $\gamma$ in the same general case as Case IV. However, they analyzed the results without estimating the critical Rayleigh number above which the convective instability can occur. We somewhat repeat the discussion here in more detail to distinguish from those presented in Cases I-III. So, in contrast to Fig. 3 of Ref. [7], we plot the wave frequency and the growth rate against the normalized horizontal wave number $k_{\perp}H$ to maintain parity with the previous results in Cases II and III. It is seen from the subplots (a) and (b) of Fig. 3 that for each profile, there exists a critical value of $k_{\perp}H$ below which both the wave frequency $\omega_{r}$ and the growth rate $\gamma$ increase and attain maximum values (This is realized as the Rayleigh number increases with $k_{\perp}H$; See Fig. 2) and above which, while $\omega_{r}$ reaches a steady state value at higher $k_{\perp}H$, the growth rate vanishes at a finite value of the same. When the parameter $\beta$ assumes a higher value or the contribution from the buoyancy force remains higher than dissipation, both the wave frequency and the growth rate are significantly enhanced (See the solid and dashed lines). However, when the contribution from the dissipation is increased by increasing a value of $\chi$, or a value of any one of $N^{2}$ and $k_{z}H$ is increased, the growth rate gets notably reduced (Compare the solid line with dotted, dash-dotted, and thick dashed line) except for $\omega_{r}$, which increases with an increase of the value of $N^{2}$ [See the solid and dash-dotted lines in subplot (a)]. We note that the presence of the BV frequency, $N^{2}$ due to the background density inhomogeneity introduces an additional term in $Ra$, which enhances the critical or minimum value of the Rayleigh number at its increasing values but reduces the growth rate. Similarly, when the vertical wave number is increased, or the wavelength of perturbation is reduced compared to the vertical scale length, the minimum value of $Ra$ is also increased but the buoyancy force still remains smaller than the dissipative force (cf. Fig. 2; $Ra_{\rm{min}}<1$ in $0.2\lesssim k_{\perp}H\lesssim 0.7$), and eventually resulting in a reduction of the growth rate.
@@
Figure 3: The real wave frequency [$\omega_{r}$, Subplot (a)] and the instability growth rate [$\gamma$, Subplot (b)], given by Eq. (23), are shown (Case IV, Sec. II.1: $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}\neq 0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $k_{z}H$. The solid, dashed, dotted, dash-dotted, and thick dashed lines, respectively, correspond to the values (i) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (ii) $\beta_{1}=0.5,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (iii) $\beta_{1}=0.25,~{}\chi_{1}=0.06,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (iv) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.41$, and $k_{z}H=0.3$, and (v) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.5$. Here, $H$ is the vertical length scale of perturbation. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are the same as for Fig. 1, i.e., $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$.
II.2 Convective flow in a bounded domain
@@
It is to be noted that the atmospheric temperature changes sharply with height, i.e., it decreases as the atmospheric height increases from a layer at $z=z_{0}$ to another layer at $z=z_{1}$. So, one can consider the tropospheric region to be bounded between two “walls” at heights $z=z_{0}$ and $z=z_{1}$ so that $z_{1}-z_{0}=H$ represents the height of the troposphere (See Fig. 4). We assume the background (or mean) temperature, $\overline{T}=T_{0}$ at $z_{0}=0$ km and $\overline{T}=T_{1}$ at $z_{1}=15$ km for the tropospheric layers. Furthermore, from the unperturbed state of the heat-energy equation, i.e., $d^{2}\overline{T}/dz^{2}=0$, we have [7]
@@
where $C_{0}$ and $C_{1}$ are constants such that $C_{0}=d\overline{T}/dz=-\Gamma$. Here, we repeat, $\Gamma$ is called the lapse rate, and $\Gamma>0$ holds for a negative temperature gradient as relevant for the troposphere [6]. Since we have assumed that $\overline{T}=T_{0}$ at $z_{0}=0$ km, we define $C_{1}=T_{0}$. Also, from the condition $\overline{T}=T_{1}$ at $z=z_{1}=H$ (See Fig. 4), we get $T_{1}=C_{0}z_{1}+C_{1}$, which gives
@@
i.e., the unperturbed temperature inhomogeneity relative to some reference temperature $T_{0}$ in a bounded domain is directly proportional to the atmospheric height $z$.
@@
Typically, the profiles of vertical perturbations depend on the boundary conditions on the upper and lower “rigid” surfaces, which are the no-slip conditions, and that the walls are maintained at constant temperatures. Thus, we choose the boundary conditions as
@@
where, in contrast to the unbounded region (Sec. II.1), we have assumed the wave amplitude of perturbations to vary with the $z$-coordinate and the wave propagates in the $xy$-plane. Substitution of Eq. (31) into Eq. (5) results in the following reduced equation.
@@
For the existence of zero-frequency waves with $\omega=0$ as in Case III of Sec. II.1, Eq. (32) gives for $N^{2}=0$ the following equation [8].
@@
where $\tilde{p}_{0}=\tilde{p}(0)$ is a constant and $m$ stands for the vertical wave number. Here, we choose a different symbol for the vertical wave number to distinguish from that for the unbounded region. Substituting the solution form of perturbation (31) into Eq. (33), we obtain the following relation between the Rayleigh number and the wave number $k_{\perp}$ [8].
@@
For a fixed vertical wave number $m$, $Ra$ takes its minimum value $Ra_{\rm{min}}=(27/4)m^{4}H^{4}$ [cf. Eq. (57)] at $k_{\perp}=m/\sqrt{2}$. From Eq. (40), $m$ takes the smallest value, $|m|=\pi/H$ at $n=0$ or $-1$. In this case, the critical values of the Rayleigh number and the horizontal wave number, above and below which the convective instability occurs, are given by [8]
@@
It follows that in the case of a bounded region, we require rather a higher value of the Rayleigh number for the instability to occur compared to the bounded region, and the growth rate can achieve a maximum value within the domain $0<k_{\perp}H\lesssim 2.2$. Similar to the Cases III and IV of Sec. II.1 with $\omega_{r}=0$ but $\gamma\neq 0$, we can obtain expressions for the instability growth rates of convective motions in a bounded region of the troposphere. From Eqs. (31) and (40), we have
@@
Also, from Eqs. (34) and (35), we have the following solutions for the transverse velocity components [8].
@@
Now, from the dispersion equation (15) of Case III, Sec. II.1, the condition for $\omega_{r}=0$ gives
@@
Next, we obtain a solution for $T^{\prime}$ from Eqs. (38) and (43) as [8]
@@
which shows that the background thermal gradient and the thermal fluctuation influence the vertical movements of tropospheric fluids in a bounded domain, which is also responsible for the instability to grow instead of damping due to the thermal diffusivity. Thus, it is pertinent to define the horizontal average of vertical heat flux as [8]
@@
For the critical mode (neutral wave) with $\gamma=0$ [See Eq. (18)] and $k^{2}=3m^{2}/2=3\pi^{2}/2H^{2}$, Eq. (51) gives [8]
@@
Next, we introduce the Nusselt number, which is the ratio between the total heat flux and the heat flux due to the thermal conductivity. Since the horizontal average of the conductivity heat flux is $\chi\Gamma$, the Nusselt number for the critical mode is given by [8]
@@
In the linear stage, the amplitude $\widetilde{T}_{0}$ is a small quantity. Even in the nonlinear regime, the amplitude of temperature perturbation is bounded by half of the temperature difference between the top and bottom boundaries, i.e., $\widetilde{T}_{0}\leq\Gamma H/2$. If we choose the maximum value of the temperature, $\widetilde{T}_{0\max}=\Gamma H/2$, the Nusselt number estimated in Eq. (53) also takes the maximum value, i.e., $Nu_{\max}=2.85$, which characterizes a kind of slug flow or laminar fluid flow. In the following, we will consider two different cases separately similar to Cases III and IV of Sec. II.1 for the unbounded region.

### v2

@@ 2.1 Convective flow in an unbounded domain
For visualization of the characteristics of the growth rate, we have plotted $\gamma$ against the normalized wave number $k_{\perp}H$ as shown in Fig. 1. It is evident that, in contrast to Cases I and II, where the buoyancy force is the only dominating force, and the perturbations grow with increasing values of $k_{\perp}H$, the growth rate has cut-offs at smaller and larger wavelengths of perturbations. First of all, from the dependency of $\gamma$ [See Eq. (16)] on the parameters, we note that the qualitative behaviors of $\gamma$ by the influences of $\beta$ and $\chi$ are the same as of $\Gamma$ and $\nu$ respectively. Thus, it is sufficient to investigate the qualitative features of $\gamma$ by the effects of the parameters $\beta$, $\chi$, $N^{2}$ (for the present case, we assume it to be zero), and the vertical wave number $k_{z}H$. Initially, for fixed values of the parameters $\beta$, $\chi$, and $k_{z}H$, and a small value of $k_{\perp}H~{}(<1)$, the contributions from both the buoyancy and dissipative forces are small, and so is the growth rate (close to zero). As the value of $k_{\perp}H$ starts increasing, the buoyancy force tends to dominate over the dissipation for which the growth rate gradually increases and reaches a maximum value until the domination is strong. As $k_{\perp}H$ further increases, the contribution from the dissipation starts increasing, and the growth rate tends to reduce and eventually vanishes at a finite value of $k_{\perp}H$ (typically $>1$). Thus, we may be conclude that in a dissipative medium of incompressible fluids under gravity in an unbounded domain of the troposphere, the perturbations can grow (leading to the convective instability) only within a finite domain of the wave number $k_{\perp}H$ due to an interplay between the buoyancy force and the dissipation. On small increasing the value of the thermal expansion coefficient $\beta$, and hence the contribution from the buoyancy force, a significant enhancement of the growth rate having a cut-off at higher $k_{\perp}H$ is seen (See the dashed line). Thus, as the wave number or the vertical length scale of perturbations deepens, the buoyancy force-driven wave motion tends to become more unstable with higher growth rates. However, such growth rates can get reduced by increasing either $\chi$ (i.e., the enhanced contribution from the dissipation) or the vertical wave number $k_{z}H$ (Since the contribution of the buoyancy force inversely varies with $k_{z}H$, it decreases with increasing $k_{z}H$; See the dotted and dash-dotted lines).
@@
Figure 1: The instability growth rate, given by Eq. (16) ($\gamma\equiv\gamma_{+}$), is shown (Case III, Sec. 2.1: $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}=0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, and $k_{z}H$ as in the legend. Here, $H$ is the vertical length scale of perturbation. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$.
@@
We note that the lower sign of $\pm$ in $\gamma$ gives a purely damped mode due to the effects of the kinematic viscosity and the thermal diffusivity, which is outside the scope of the present study. Thus, for the upper sign, the instability condition $(\gamma>0)$ gives
@@
From Eq. (25), we find that for a given vertical scale height $H$, the term proportional to $\beta$, associated with the buoyancy force due to thermal expansion and background density inhomogeneity, reduces the critical (or minimum) value of the Rayleigh number [cf. Eq. (20)] above which the convective instability can occur. Typically, for $H\sim 800$ m, $N^{2}\sim-0.003/\rm{s}^{2}$, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=260$ K, $\Gamma=0.0065$ K/m, $\beta=0.001$/K, $\chi=3000~{}\rm{m^{2}/s}$, and $\nu=2000~{}\rm{m^{2}/s}$, relevant for tropospheric fluids (Kaladze and Misra, 2023), the variation of the critical values of the Rayleigh number against the normalized wave number $k_{\perp}H$ are shown in Fig. 2. We see that after the sharp peak at a point close to the singular point $k_{\perp}H=0$, the critical value of the Rayleigh number decreases, achieving a minimum ($\approx 0.18$) at $k_{\perp}H\approx 0.35$, and then tends to increase slowly with increasing values of $k_{\perp}H$. Thus, given a small but finite value of the wave number $k_{\perp}$, a larger vertical length scale of perturbations $H$ results in higher values of $Ra_{\rm{min}}$ and the gravitational force to become more dominant over the pressure gradient and viscous forces, which can lead to stronger convective instability. Mathematically, the minimum value of the Rayleigh number can be exceedingly high at $k_{\perp}H\lesssim 0.1$, i.e., when the wavelength of horizontal perturbations greatly exceeds the vertical scale size $H$. However, this limiting case may not be relevant to unbounded vertical domains (with infinite extent) of the troposphere. We also note that the critical Rayleigh number assumes values less than unity in the interval $0.2\lesssim k_{\perp}H\lesssim 0.7$, corresponding to a regime where the dissipative force is stronger than the buoyancy force. In this case, for the instability to occur, the Rayleigh number can be smaller than unity even when it exceeds the critical value. Such a small value of $Ra<1$ and the corresponding regime of $k_{\perp}H$ may not be admissible as we require the buoyancy force to be dominant over the dissipative force for stronger instability. Thus, the convective wave motion of tropospheric viscous fluids under gravity becomes more unstable when the wavelengths of horizontal perturbations remain smaller compared to the vertical scale size (i.e., $k_{\perp}^{-1}<H$). We note that the critical value of the Rayleigh number does not alter significantly with a small change of the values of the parameters $\beta$ (or $\Gamma$), $\chi$ (or $\nu$), and $N^{2}$. Also, at $Ra=Ra_{\rm{min}}$, the growth rate vanishes, and one can have the convective motion of neutral waves for $N^{2}<0$ with the wave frequency, given by,
@@
Figure 2: The critical value (or minimum value) of the Rayleigh number, given by Eq. (25), is plotted against the perpendicular component of the normalized wave number $k_{\perp}H$. Apart from the sharp peak at a point close to the singular point $k_{\perp}H=0$, the critical value tends to increase slowly with increasing values of $k_{\perp}H$. The fixed parameter values are $\beta=0.001$/K, $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}=0.04$, $k_{z}H=0.5$, and $N^{2}=-0.003/\rm{s}^{2}$. The minimum value of the Rayleigh number is achieved as $Ra_{\rm{min}}\approx 0.18$ at $k_{\perp}H\approx 0.35$.
@@
In particular, for $N^{2}=0$, i.e., when we ignore the density inhomogeneity, the condition (24) reduces to the same condition as Eq. (17).
@@
Figure 3 shows the profiles of the real wave frequency $\omega_{r}$ [Subplot (a)] and the growth rate of instability $\gamma$ [Subplot (b)], given by Eq. (23), for different values of the parameters $\beta$, $\chi$, $N^{2}$, and $k_{z}$. The influences of the other parameters on the growth rate are not shown for the reason mentioned before in Case III. The authors of Ref. (Kaladze and Misra, 2024) discussed the characteristics of $\omega_{r}$ and $\gamma$ in the same general case as Case IV. However, they analyzed the results without estimating the critical Rayleigh number above which the convective instability can occur. We somewhat repeat the discussion here in more detail to distinguish from those presented in Cases I-III. So, in contrast to Fig. 3 of Ref. (Kaladze and Misra, 2024), we plot the wave frequency and the growth rate against the normalized horizontal wave number $k_{\perp}H$ to maintain parity with the previous results in Cases II and III. It is seen from the subplots (a) and (b) of Fig. 3 that for each profile, there exists a critical value of $k_{\perp}H$ below which both the wave frequency $\omega_{r}$ and the growth rate $\gamma$ increase and attain maximum values (This occurs as the Rayleigh number increases with $k_{\perp}H$; See Fig. 2) and above which, while $\omega_{r}$ reaches a steady state value at higher $k_{\perp}H$, the growth rate vanishes at a finite value of the same. When the parameter $\beta$ assumes a higher value or the contribution from the buoyancy force remains higher than dissipation, both the wave frequency and the growth rate are significantly enhanced (See the solid and dashed lines). However, when the contribution from the dissipation gets increased by increasing a value of $\chi$, or a value of any one of $N^{2}$ and $k_{z}H$ is increased, the growth rate gets notably reduced (Compare the solid line with dotted, dash-dotted, and thick dashed line) except for $\omega_{r}$, which increases with an increase of the value of $N^{2}$ [See the solid and dash-dotted lines in Subplot (a)]. We note that the presence of the BV frequency, $N^{2}$, due to the background density inhomogeneity introduces an additional term in $Ra$, which enhances the critical or minimum value of the Rayleigh number at its increasing values but reduces the growth rate. Similarly, when the vertical wave number gets increased, or the wavelength of perturbation is reduced compared to the vertical scale length, the minimum value of $Ra$ is also increased, but the buoyancy force remains smaller than the dissipative force (cf. Fig. 2; $Ra_{\rm{min}}<1$ in $0.2\lesssim k_{\perp}H\lesssim 0.7$), and eventually resulting in a reduction of the growth rate.
@@
Figure 3: The real wave frequency [$\omega_{r}$, Subplot (a)] and the instability growth rate [$\gamma$, Subplot (b)], given by Eq. (23), are shown (Case IV, Sec. 2.1: $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}\neq 0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $k_{z}H$. The solid, dashed, dotted, dash-dotted, and thick dashed lines, respectively, correspond to the values (i) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (ii) $\beta_{1}=0.5,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (iii) $\beta_{1}=0.25,~{}\chi_{1}=0.06,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (iv) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.41$, and $k_{z}H=0.3$, and (v) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.5$. Here, $H$ is the vertical length scale of perturbation. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are the same as for Fig. 1, i.e., $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$.
2.2 Convective flow in a bounded domain
@@
It is to be noted that the atmospheric temperature changes sharply with height, i.e., it decreases as the atmospheric height increases from a layer at $z=z_{0}$ to another layer at $z=z_{1}$. So, one can consider the tropospheric region to be bounded between two “walls" at heights $z=z_{0}$ and $z=z_{1}$ so that $z_{1}-z_{0}=H$ represents the height of the troposphere (See Fig. 4). We assume the background (or mean) temperature, $\overline{T}=T_{0}$ at $z_{0}=0$ km and $\overline{T}=T_{1}$ at $z_{1}=15$ km for the tropospheric layers. Furthermore, from the unperturbed state of the heat-energy equation, i.e., $d^{2}\overline{T}/dz^{2}=0$, we have (Kaladze and Misra, 2024)
@@
where $C_{0}$ and $C_{1}$ are constants such that $C_{0}=d\overline{T}/dz=-\Gamma$. Here, we repeat, $\Gamma$ is called the lapse rate, and $\Gamma>0$ holds for a negative temperature gradient as relevant for the troposphere (Kaladze and Misra, 2023). Since we have assumed that $\overline{T}=T_{0}$ at $z_{0}=0$ km, we define $C_{1}=T_{0}$. Also, from the condition $\overline{T}=T_{1}$ at $z=z_{1}=H$ (See Fig. 4), we get $T_{1}=C_{0}z_{1}+C_{1}$, which gives
@@
i.e., the unperturbed temperature inhomogeneity relative to some reference temperature $T_{0}$ in a bounded domain is directly proportional to the atmospheric height $z$. It follows that the values of $N^{2}$ will differ in unbounded and bounded regions of atmospheric fluids.
@@
Typically, the profiles of vertical perturbations depend on the boundary conditions on the upper and lower “rigid" surfaces, which are the no-slip conditions, and that the walls are maintained at constant temperatures. Thus, we choose the boundary conditions as
@@
where, in contrast to the unbounded region (Sec. 2.1), we have assumed the wave amplitude of perturbations to vary with the $z$-coordinate and the wave propagates in the $xy$-plane. Substitution of Eq. (31) into Eq. (5) results in the following reduced equation.
@@
For the existence of zero-frequency waves with $\omega=0$ as in Case III of Sec. 2.1, Eq. (32) gives for $N^{2}=0$ the following equation (Satoh, 2004).
@@
where $\tilde{p}_{0}=\tilde{p}(0)$ is a constant and $m$ stands for the vertical wave number. Here, we choose a different symbol for the vertical wave number to distinguish from that for the unbounded region. Substituting the solution form of perturbation (31) into Eq. (33), we obtain the following relation between the Rayleigh number and the wave number $k_{\perp}$ (Satoh, 2004).
@@
For a fixed vertical wave number $m$, $Ra$ takes its minimum value $Ra_{\rm{min}}=(27/4)m^{4}H^{4}$ [cf. Eq. (57)] at $k_{\perp}=m/\sqrt{2}$. From Eq. (40), $m$ takes the smallest value, $|m|=\pi/H$ at $n=0$ or $-1$. In this case, the critical values of the Rayleigh number and the horizontal wave number, above and below which the convective instability occurs, are given by (Satoh, 2004)
@@
It follows that in the case of a bounded region, we require a higher value of the Rayleigh number for the instability to occur compared to the bounded region, and the growth rate can achieve a maximum value within the domain $0<k_{\perp}H\lesssim 2.2$. Similar to cases III and IV of Sec. 2.1 with $\omega_{r}=0$ but $\gamma\neq 0$, we can obtain expressions for the instability growth rates of convective motions in a bounded region of the troposphere. From Eqs. (31) and (40), we have
@@
Also, from Eqs. (34) and (35), we have the following solutions for the transverse velocity components (Satoh, 2004).
@@
Now, from the dispersion equation (15) of Case III, Sec. 2.1, the condition for $\omega_{r}=0$ gives
@@
Next, we obtain a solution for $T^{\prime}$ from Eqs. (38) and (43) as (Satoh, 2004)
@@
which shows that the background thermal gradient and the thermal fluctuation influence the vertical movements of tropospheric fluids in a bounded domain, which is also responsible for the instability to grow instead of damping due to the thermal diffusivity. Thus, it is pertinent to define the horizontal average of vertical heat flux as (Satoh, 2004)
@@
For the critical mode (neutral wave) with $\gamma=0$ [See Eq. (18)] and $k^{2}=3m^{2}/2=3\pi^{2}/2H^{2}$, Eq. (51) gives (Satoh, 2004)
@@
Next, we introduce the Nusselt number, which is the ratio between the total heat flux and the heat flux due to the thermal conductivity. Since the horizontal average of the conductivity heat flux is $\chi\Gamma$, the Nusselt number for the critical mode is given by (Satoh, 2004)
@@
The amplitude $\widetilde{T}_{0}$ in the linear regime is small. However, in the nonlinear regime, the amplitude of temperature perturbation is bounded by half of the temperature difference between the top and bottom boundaries, i.e., $\widetilde{T}_{0}\leq\Gamma H/2$. If we choose the maximum value of the temperature, $\widetilde{T}_{0\max}=\Gamma H/2$, the Nusselt number estimated in Eq. (53) also takes the maximum value, i.e., $Nu_{\max}=2.85$, which characterizes a kind of slug flow or laminar fluid flow. In the following, we will consider two different cases separately, similar to Cases III and IV of Sec. 2.1 for the unbounded region.